please help me solve this problem. a chemist has one solution that is 15% acid and another solution that is 10% acid. how much of the 15% solution is needed to make 80 L of the solution that is 12% acid
Answer by ewatrrr(24785)


Hi,
Let x represent the amount of the 15% solution needed.
Then (80L-x) would be the amount of 10% solution
.15x + .10(80L-x) = .12*80L
solve for x
.05x = .02*80L
x = 32L of the 15% solution is needed
checking our answer
.15*32L + .10*48 = 4.8 + 4.8 = 9.6